Suppose the height, h, in feet, of a trampolinist above the ground during one bounce is modeled by the quadratic function h(t) = -16t^2 + 42t + 3.75 . For what period of time is the trampolinist at least 22 ft above the ground? Round your answers to the nearest hundredth.
Answer by TimothyLamb(4379)

22 = -16t^2 + 42t + 3.75
-16t^2 + 42t + 3.75 - 22 = 0
-16t^2 + 42t - 18.25 = 0
---
the above quadratic equation is in standard form, with a=-16, b=42 and c=-18.25
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
-16 42 -18.25
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
t = 0.549590274
t = 2.07540973
---
answer:
he is at least 22 ft above the ground from t=0.55 and t=2.08 seconds
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---